# AdaSplash: Adaptive Sparse Flash Attention

### v1

@@ Abstract
The computational cost of softmax-based attention in transformers limits their applicability to long-context tasks. Adaptive sparsity, of which $\alpha$-entmax attention is an example, offers a flexible data-dependent alternative, but existing implementations are inefficient and do not leverage the sparsity to obtain runtime and memory gains. In this work, we propose AdaSplash, which combines the efficiency of GPU-optimized algorithms with the sparsity benefits of $\alpha$-entmax. We first introduce a hybrid Halley-bisection algorithm, resulting in a 7-fold reduction in the number of iterations needed to compute the $\alpha$-entmax transformation. Then, we implement custom Triton kernels to efficiently handle adaptive sparsity. Experiments with RoBERTa and ModernBERT for text classification and single-vector retrieval, along with GPT-2 for language modeling, show that our method achieves substantial improvements in runtime and memory efficiency compared to existing $\alpha$-entmax implementations. It approaches—and in some cases surpasses—the efficiency of highly optimized softmax implementations like FlashAttention-2, enabling long-context training while maintaining strong task performance.
Sparsity, Flash Attention
@@ 1 Introduction
Central to the success of transformers (Vaswani et al., 2017) lies the attention mechanism, where each token in a sequence attends directly to every other token. Attention probabilities are computed through the softmax transformation, which always assigns a nonzero probability to every token. However, for long context inputs, the accumulation of small probabilities can lead to dispersion (Veličković et al., 2024). In fact, previous research shows that attention probabilities tend to peak around a small number of tokens (Voita et al., 2019; Treviso et al., 2022), which suggests that model performance and computational efficiency can be increased by leveraging attention sparsity. This has motivated methods that predefine sparse masks (Beltagy et al., 2020b; Zaheer et al., 2020), rely on clustering-based strategies (Kitaev et al., 2020), or low-rank approximate attention (Choromanski et al., 2021; Peng et al., 2021; Xiong et al., 2021; Chen et al., 2021). Some of these techniques show the potential of sparsity to mitigate memory and computation bottlenecks, but they often require architectural modifications or crude approximations, limiting their flexibility and generality.
@@
This paper addresses this problem by providing new algorithms and implementations to improve the computational efficiency of the family of $\alpha$-entmax activations. Our main contributions include a faster and GPU-friendly algorithm for calculating $\alpha$-entmax, alongside a Triton kernel (Tillet et al., 2019) for computing entmax-based attention, which we call AdaSplash. In particular, AdaSplash advances the goal of supporting training of adaptively sparse models with longer context lengths, as shown in Figure 1. We demonstrate the potential and scalability of our approach through experiments with synthetic data and with several natural language processing benchmarks for encoder-only and decoder-only models, achieving substantial improvements over previous $\alpha$-entmax implementations and approaching (sometimes surpassing) the efficiency of softmax-based attention with FlashAttention-2, with strong performance on downstream tasks. [^1] Code at: https://github.com/deep-spin/adasplash
@@ 2.1 Hardware Performance
Modern GPUs, such as the Nvidia H100, are designed for efficient parallel computation using a hierarchical memory architecture, with high-bandwidth memory (HBM) providing large capacity but slower access compared to the smaller, faster on-chip SRAM. Efficient use of SRAM is critical to minimize the memory bottlenecks caused by frequent HBM accesses. GPUs execute operations (kernels) via thousands of threads organized into blocks and warps, where data is loaded from HBM into SRAM for computation before being written back. Kernel fusion is a key optimization strategy that combines multiple operations into a single kernel, reducing intermediate HBM accesses by directly computing and storing final results. While compilers like torch.compile can automate fusion for simple operations (Ansel et al., 2024), complex tasks such as attention mechanisms require custom strategies to reorder operations and optimize memory usage effectively. Our method leverages this GPU memory organization by implementing block-wise computations, recomputation strategies, and kernel fusion specifically tailored for sparse attention, as detailed in §3.2.1 and §3.2.2.
@@ 2.4 Sparse Attention
The original softmax-based attention is dense, i.e., it puts some probability mass on all tokens—not only a computational disadvantage, but also making interpretation and generalization harder (Voita et al., 2019; Treviso et al., 2022; Veličković et al., 2024). An alternative to softmax is the $\bm{\alpha}$-entmax transformation (Peters et al., 2019), which is differentiable and leads to sparse outputs:
@@
where $[\cdot]_{+}$ is the ReLU function, and $\tau\in\mathbb{R}$ is a normalizing constant to ensure the output is a valid probability distribution. Importantly, entries with score $s_{i}\leq\frac{\tau}{\alpha-1}$ get exactly zero probability. In the limit $\alpha\rightarrow 1$, $\alpha$-entmax recovers the softmax function, while for any value of $\alpha>1$ this transformation returns increasingly sparser probability vectors. When $\alpha=2$, we recover the sparsemax transformation (Martins & Astudillo, 2016). However, in contrast to fixed sparse patterns, such as windowed sparse attention (Child et al., 2019; Beltagy et al., 2020b) and block-sparse variants (Zaheer et al., 2020; Dao et al., 2022), $\alpha$-entmax’s sparsity patterns are dynamic and hence difficult to exploit in order to reduce the quadratic burden of self-attention because we still need to materialize $\bm{S}=\bm{Q}\bm{K}^{\top}$ before applying the transformation.
@@ 3.2.1 Forward Pass
The key idea involves splitting the inputs $\bm{Q},\bm{K},\bm{V}$ into smaller blocks, and then computing attention block by block. We start by loading only $\bm{Q}$ and $\bm{K}$ from the slower HBM to the faster SRAM to compute $\bm{\tau}\in\mathbb{R}^{n}$ using the Halley-bisection algorithm (Alg. 1). In order to use the aforementioned algorithm, we need to accumulate three values: $f(\bm{\tau})$, $f^{\prime}(\bm{\tau})$, $f^{\prime\prime}(\bm{\tau})$. Since $f$, as well at its derivatives, is additive over its inputs, their computation can also be computed in blocks. Let $B_{r}$ and $B_{c}$ be the row and column block sizes, respectively, and define $T_{r}=\lceil n/B_{r}\rceil$ and $T_{c}=\lceil n/B_{c}\rceil$. Divide $\bm{Q}$ into $\bm{Q}_{1},...,\bm{Q}_{T_{r}}$ blocks, and $\bm{K}$ into $\bm{K}_{1},...,\bm{K}_{T_{c}}$ blocks. Then, $f(\bm{\tau})$ can be computed as:
@@ 4.2 Performance on Real Tasks
We fine-tuned a pretrained RoBERTa model (Liu et al., 2019) on the ECtHR (Chalkidis et al., 2019, 2021) dataset while progressively increasing the sequence length up to 8192 tokens. Positional embeddings were extended by repetition, following the approach of Beltagy et al. (2020b). As a baseline, we fine-tuned the model using standard softmax-based attention. For $\alpha$-entmax attention, we linearly increased the $\alpha$ from 1.0 to 1.5 during training to ensure smooth convergence. The results, summarized in Table 2, show a consistent improvement in model performance with longer context lengths. Notably, despite the base model being pretrained with standard attention, $\alpha$-entmax attention was capable of effectively learning the task, achieving a slightly higher micro $F_{1}$ score than the model fine-tuned with standard attention up to a sequence length of 4096 tokens.
@@
Figure 4: Ratio of non-zero attention scores for GPT-2 ($\alpha=1.5$).
@@ 5 Related Works
Efficient sparse attention mechanisms have been widely studied to reduce the quadratic cost of transformers. The Sparse Transformer (Child et al., 2019) introduces a fixed windowed attention that can be efficiently computed using CUDA kernels, a strategy also adopted by Longformer (Beltagy et al., 2020a). However, data-dependent sparse attention methods, such as Reformer (Kitaev et al., 2020) and Routing Transformer (Roy et al., 2021), still rely on softmax and do not leverage the sparsity of attention weights. Adaptively sparse transformers (Correia et al., 2019) offer an alternative approach where attention heads can learn $\alpha$ dynamically, improving interpretability but without leveraging sparsity for efficiency. SparseFinder (Treviso et al., 2022) aims to address this issue by predicting the sparsity pattern of entmax attention a priori; however, it does not scale efficiently for long sequences.
@@
Recent works have explored optimizing attention mechanisms with hardware-aware implementations. Flex Attention (Dong et al., 2024) provides an API for efficient attention computation, though they remain tied to softmax-based transformations and do not support more complex operations such as those considered in our work. Closely related to our approach, FlashAttention-1 and 2 (Dao et al., 2022; Dao, 2024) optimize softmax-based attention using tiling and recomputation techniques implemented in CUDA. While FlashAttention-1 includes a sparse block variant, its sparsity pattern must be predefined, limiting adaptability. In this work, we compare our method, AdaSplash, with FlashAttention-2 and demonstrate that our approach can outperform both its CUDA and Triton implementations at high input sparsity levels. Similarly, Sparse Flash Attention (Pagliardini et al., 2023) extends FlashAttention-1 with a sparse variant that reduces computational cost by either dropping queries and keys per head or grouping them using a hash-based bucketing approach. However, despite its efficiency improvements, it relies on slow sorting operations and is constrained to causal attention, making its sparsity a by-product of bucketing rather than an inherently adaptive feature, as in our case.
@@ Acknowledgments
We thank António Farinhas, Ben Peters, Duarte Alves, Giuseppe Attanasio, Saul Santos, Edoardo Ponti, and Vlad Niculae for their helpful and constructive feedback. This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ B.3 Long Document Classification
The European Court of Human Rights (ECtHR) dataset comprises legal cases from the European Court of Human Rights, each associated with specific articles of the Convention on Human Rights allegedly violated. For this task, we fine-tuned the RoBERTa base model (Liu et al., 2019) with a classification head. Since this is a multi-label classification task, we used a binary cross-entropy loss. To accommodate longer contexts, we followed the approach proposed by (Beltagy et al., 2020b), repeating the 512 position embeddings until the target context size was reached. We used the AdamW optimizer for training. For hyperparameters, we follow the recipe of Dai et al. (2022). For the attention mechanism, bfloat16 precision was used.
Table 6: Runtime per epoch (hh:mm:ss) and peak memory usage (GB) for different sequence lengths on RoBERTa (FA2) and RoBERTa ($\alpha=1.5$) with AdaSplash and Torch’s bisection for the Long Document Classification task. In cases where the full batch could not fit in memory, gradient accumulation was used. Memory values represent the effective peak memory required to process a batch of 16, given the gradient accumulation steps.
Runtime (hh:mm:ss) Memory (GB)
Model 512 1024 2048 4096 8192 512 1024 2048 4096 8192
RoBERTa (FA2) 2:39 5:00 9:35 18:36 35:51 6.75 11.43 20.35 37.49 75.00
RoBERTa ($\alpha=1.5$) 2:45 5:20 10:24 19:54 38:08 6.75 11.45 20.38 39.17 79.88
     w/ Torch Bisect 4:51 8:44 22:48 1:11:53 4:12:34 7.75 16.92 44.06 142.76 508.16
@@ Appendix C Computational Details
Experiments on masked language modeling, text classification, GLUE tasks and BIER tasks were carried on Nvidia RTX A6000 GPUs with 48GB VRAM. Experiments with GPT-2 and the efficiency benchmark in Figures 1 and 3 were carried on a single Nvidia H100 Tensor Core GPU with 80GB VRAM. The runtime experiments with ModernBERT were carried on a single A6000 GPU.

### v2

@@ Abstract
The computational cost of softmax-based attention in transformers limits their applicability to long-context tasks. Adaptive sparsity, of which $\alpha$-entmax attention is an example, offers a flexible data-dependent alternative, but existing implementations are inefficient and do not leverage the sparsity to obtain runtime and memory gains. In this work, we propose AdaSplash, which combines the efficiency of GPU-optimized algorithms with the sparsity benefits of $\alpha$-entmax. We first introduce a hybrid Halley-bisection algorithm, resulting in a 7-fold reduction in the number of iterations needed to compute the $\alpha$-entmax transformation. Then, we implement custom Triton kernels to efficiently handle adaptive sparsity. Experiments with RoBERTa and ModernBERT for text classification and single-vector retrieval, along with GPT-2 for language modeling, show that our method achieves substantial improvements in runtime and memory efficiency compared to existing $\alpha$-entmax implementations. It approaches—and in some cases surpasses—the efficiency of highly optimized softmax implementations like FlashAttention-2, enabling long-context training while maintaining strong task performance. [^1] Code: https://github.com/deep-spin/adasplash
Machine Learning, ICML
@@ 1 Introduction
Central to the success of transformers (Vaswani et al., 2017) lies the attention mechanism, where each token in a sequence attends directly to every other token. Attention probabilities are computed through the softmax transformation, which always assigns a nonzero probability to every token. However, for long context inputs, the accumulation of small probabilities can lead to dispersion (Veličković et al., 2025). In fact, previous research shows that attention probabilities tend to peak around a small number of tokens (Voita et al., 2019; Treviso et al., 2022), which suggests that model performance and computational efficiency can be increased by leveraging attention sparsity. This has motivated methods that predefine sparse masks (Beltagy et al., 2020; Zaheer et al., 2020b), rely on clustering-based strategies (Kitaev et al., 2020), or low-rank approximate attention (Choromanski et al., 2021; Peng et al., 2021; Xiong et al., 2021; Chen et al., 2021). Some of these techniques show the potential of sparsity to mitigate memory and computation bottlenecks, but they often require architectural modifications or crude approximations, limiting their flexibility and generality.
@@
This paper addresses this problem by providing new algorithms and implementations to improve the computational efficiency of the family of $\alpha$-entmax activations. Our main contributions include a faster and GPU-friendly algorithm for calculating $\alpha$-entmax, alongside a Triton kernel (Tillet et al., 2019) for computing entmax-based attention, which we call AdaSplash. In particular, AdaSplash advances the goal of supporting training of adaptively sparse models with longer context lengths, as shown in Figure 1. We demonstrate the potential and scalability of our approach through experiments with synthetic data and with several natural language processing benchmarks for encoder-only and decoder-only models, achieving substantial improvements over previous $\alpha$-entmax implementations and approaching (sometimes surpassing) the efficiency of softmax-based attention with FlashAttention-2, with strong performance on downstream tasks.
@@ 2.1 Hardware Performance
Modern GPUs, such as the Nvidia H100, are designed for efficient parallel computation using a hierarchical memory architecture, with high-bandwidth memory (HBM) providing large capacity but slower access compared to the smaller, faster on-chip SRAM. Efficient use of SRAM is critical to minimize the memory bottlenecks caused by frequent HBM accesses. GPUs execute operations (kernels) via thousands of threads organized into thread blocks, where data is loaded from HBM into SRAM for computation before being written back. Kernel fusion is a key optimization strategy that combines multiple operations into a single kernel, reducing intermediate HBM accesses by directly computing and storing final results. While compilers like torch.compile can automate fusion for simple operations (Ansel et al., 2024), complex tasks such as attention mechanisms require custom strategies to reorder operations and optimize memory usage effectively. Our method leverages this GPU memory organization by implementing block-wise computations, recomputation strategies, and kernel fusion specifically tailored for sparse attention, as detailed in §3.2.1 and §3.2.2.
@@ 2.4 Sparse Attention
The original softmax-based attention is dense, i.e., it puts some probability mass on all tokens—not only a computational disadvantage, but also making interpretation and generalization harder (Voita et al., 2019; Treviso et al., 2022; Veličković et al., 2025). An alternative to softmax is the $\bm{\alpha}$-entmax transformation (Peters et al., 2019), which is differentiable and leads to sparse outputs:
@@
where $[\cdot]_{+}$ is the ReLU function, and $\tau\in\mathbb{R}$ is a normalizing constant to ensure the output is a valid probability distribution. Importantly, entries with score $s_{i}\leq\frac{\tau}{\alpha-1}$ get exactly zero probability. In the limit $\alpha\rightarrow 1$, $\alpha$-entmax recovers the softmax function, while for any value of $\alpha>1$ this transformation returns increasingly sparser probability vectors. When $\alpha=2$, we recover the sparsemax transformation (Martins & Astudillo, 2016). However, in contrast to fixed sparse patterns, such as windowed sparse attention (Child et al., 2019; Beltagy et al., 2020) and block-sparse variants (Zaheer et al., 2020b; Dao et al., 2022), $\alpha$-entmax’s sparsity patterns are dynamic and hence difficult to exploit in order to reduce the quadratic burden of self-attention because we still need to materialize $\bm{S}=\bm{Q}\bm{K}^{\top}$ before applying the transformation.
@@ 3.2.1 Forward Pass
The key idea involves splitting the inputs $\bm{Q},\bm{K},\bm{V}$ into smaller blocks, and then computing attention block by block. We start by loading only $\bm{Q}$ and $\bm{K}$ from the slower HBM to the faster SRAM to compute $\bm{\tau}\in\mathbb{R}^{n}$ using the Halley-bisection algorithm (Alg. 1). In order to use the aforementioned algorithm, we need to accumulate three values: $f(\bm{\tau})$, $f^{\prime}(\bm{\tau})$, $f^{\prime\prime}(\bm{\tau})$. Since $f$, as well as its derivatives, is additive over its inputs, their computation can also be computed in blocks. Let $B_{r}$ and $B_{c}$ be the row and column block sizes, respectively, and define $T_{r}=\lceil n/B_{r}\rceil$ and $T_{c}=\lceil n/B_{c}\rceil$. Divide $\bm{Q}$ into $\bm{Q}_{1},...,\bm{Q}_{T_{r}}$ blocks, and $\bm{K}$ into $\bm{K}_{1},...,\bm{K}_{T_{c}}$ blocks. Then, $f(\bm{\tau})$ can be computed as:
@@ 4.2 Performance on Real Tasks
Table 3: Runtime per epoch (hh:mm:ss) and peak memory usage (GB) for long document classification with different sequence lengths. In cases where the full batch could not fit in memory, gradient accumulation was used. Memory values represent the effective peak memory required to process a batch of 16 samples.
Runtime (hh:mm:ss) Sequence Length
Model 512 1024 2048 4096 8192
RoBERTa 2:39 5:00 9:35 18:36 35:51
RoBERTa ($\alpha=1.5$) 2:45 5:20 10:24 19:54 38:08
w/ Torch Bisect 4:51 8:44 22:48 1:11:53 4:12:34
Memory (GB) Sequence Length
512 1024 2048 4096 8192
RoBERTa 6.75 11.43 20.35 37.49 75.00
RoBERTa ($\alpha=1.5$) 6.75 11.45 20.38 39.17 79.88
w/ Torch Bisect 7.75 16.92 44.06 142.76 508.16
@@
We fine-tuned a pretrained RoBERTa model (Liu et al., 2019) on the ECtHR (Chalkidis et al., 2019, 2021) dataset while progressively increasing the sequence length up to 8192 tokens. Positional embeddings were extended by repetition, following the approach of Beltagy et al. (2020). As a baseline, we fine-tuned the model using standard softmax-based attention. For $\alpha$-entmax attention, we linearly increased the $\alpha$ from 1.0 to 1.5 during training to ensure smooth convergence. The results, summarized in Table 2, show a consistent improvement in model performance with longer context lengths. Notably, despite the base model being pretrained with standard attention, $\alpha$-entmax attention was capable of effectively learning the task, achieving a slightly higher micro $F_{1}$ score than the model fine-tuned with standard attention up to a sequence length of 4096 tokens.
Table 3 compares the runtime per epoch and peak memory usage for different sequence lengths on the long document classification task. We report results for RoBERTa with FlashAttention-2 ($\alpha=1$), RoBERTa with AdaSplash ($\alpha=1.5$), and RoBERTa using Torch’s bisection-based implementation. AdaSplash enables scalable training with $\alpha$-entmax attention. Prior to this, implementations had to resort to Torch’s bisection, which leads to both extremely slow runtimes or even out-of-memory problems, rendering it infeasible for most realistic training setups. In contrast, our method brings the cost of $\alpha$-entmax attention down to the level of existing dense attention implementations, as both runtime and memory usage with AdaSplash remain well aligned with those of FlashAttention-2.
@@ 5 Related Works
Efficient sparse attention mechanisms have been widely studied to reduce the quadratic cost of transformers. The Sparse Transformer (Child et al., 2019) introduces a fixed windowed attention that can be efficiently computed using CUDA kernels, a strategy also adopted by Longformer (Beltagy et al., 2020), and BigBird (Zaheer et al., 2020a). However, data-dependent sparse attention methods, such as Reformer (Kitaev et al., 2020) and Routing Transformer (Roy et al., 2021), aimi to approximate softmax in return for efficiency, not leveraging the sparsity of attention weights. Other methods, such as Top-$k$ attention (Gupta et al., 2021) and NSA (Yuan et al., 2025), provide sparsity but require a fixed, non-adaptable budget. In contrast, $\alpha$-entmax attention provides natural, input-dependent sparsity patterns with an exact and differentiable transformation that generalizes softmax, making it more flexible for modeling attention distributions. Adaptively sparse transformers (Correia et al., 2019) uses $\alpha$-entmax attention where attention heads can learn $\alpha$ dynamically, improving interpretability but without leveraging sparsity for efficiency. SparseFinder (Treviso et al., 2022) aims to address efficiency issues by predicting the sparsity pattern of entmax attention a priori; however, it does not scale efficiently for long sequences.
Figure 4: Ratio of non-zero attention scores for GPT-2 ($\alpha=1.5$).
@@
Recent works have explored optimizing attention mechanisms with hardware-aware implementations. Flex Attention (Dong et al., 2024) provides an API for efficient attention computation, though they remain tied to softmax-based transformations and do not support more complex operations such as those considered in our work. Closely related to our approach, FlashAttention-1 and 2 (Dao et al., 2022; Dao, 2024) optimize softmax-based attention using tiling and recomputation techniques implemented in CUDA. While FlashAttention includes a sparse block variant, its sparsity pattern must be predefined, limiting adaptability. In this work, we compare our method, AdaSplash, with FlashAttention-2 and demonstrate that our approach can outperform both its CUDA and Triton implementations at high input sparsity levels. Similarly, Sparse Flash Attention (Pagliardini et al., 2023) extends FlashAttention-1 with a sparse variant that reduces computational cost by either dropping queries and keys per head or grouping them using a hash-based bucketing approach. However, despite its efficiency improvements, it relies on slow sorting operations and is constrained to causal attention, making its sparsity a by-product of bucketing rather than an inherently adaptive feature, as in our case.
@@ Acknowledgments
We thank Vlad Niculae for his insightful and constructive comments throughout this work. We also thank the SARDINE Lab members for reviewing this paper and providing helpful feedback. This work was supported by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for ResponsibleAI), by the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.
@@ B.3 Long Document Classification
The European Court of Human Rights (ECtHR) dataset comprises legal cases from the European Court of Human Rights, each associated with specific articles of the Convention on Human Rights allegedly violated. For this task, we fine-tuned the RoBERTa base model (Liu et al., 2019) with a classification head. Since this is a multi-label classification task, we used a binary cross-entropy loss. To accommodate longer contexts, we followed the approach proposed by (Beltagy et al., 2020), repeating the 512 position embeddings until the target context size was reached. We used the AdamW optimizer for training. For hyperparameters, we follow the recipe of Dai et al. (2022). For the attention mechanism, bfloat16 precision was used.
@@ Appendix C Computational Details
Experiments on masked language modeling, text classification, GLUE tasks and BIER tasks were carried on Nvidia RTX A6000 GPUs with 48GB VRAM. Experiments with GPT-2 and the efficiency benchmark in Figures 1 and 3 were carried on a single Nvidia H100 GPU (80GB). The runtime experiments with ModernBERT were carried on a single A6000 GPU.